OpenStudy (anonymous):

How many grams of water formed by reacting 2gm of H2 with 2gm of O2

OpenStudy (anonymous):

@Abmon98 .... Bro any idea about this

OpenStudy (abmon98):

H2+1/2O2--->H2O 2/2=1 moles 2/32=0.0625 use the limiting reagent which is oxygen, 0.0625*18=1.125 grams although iam not sure

OpenStudy (anonymous):

but the options are...A) 2gms B) 2.25gms C)4gms D) 4.5gms

OpenStudy (abmon98):

iam sorru i didnt multiply the number of moles of oxygen by 2 because the ration 1/2:1

OpenStudy (matt101):

The answer is B. It's true that oxygen is the limiting reagent, but you need to take the stoichiometric ratios into account when finding the mass produced. 1 H2O/ ½ O2 = x H2O / 0.0625 O2 x = 0.125 So you have 0.125 moles of H2O produced, which is 18 x 0.125 = 2.25 g.

OpenStudy (abmon98):

ratio*

OpenStudy (anonymous):

@Abmon and @matt101......Thanks man...